# InstaGen: Enhancing Object Detection by Training on Synthetic Dataset

### 20 Feb 2024 v2

@@ Abstract
In this paper, we present a novel paradigm to enhance the ability of object detector, e.g., expanding categories or improving detection performance, by training on synthetic dataset generated from diffusion models. Specifically, we integrate an instance-level grounding head into a pre-trained, generative diffusion model, to augment it with the ability of localising arbitrary instances in the generated images. The grounding head is trained to align the text embedding of category names with the regional visual feature of the diffusion model, using supervision from an off-the-shelf object detector, and a novel self-training scheme on (novel) categories not covered by the detector. We conduct thorough experiments to show that, this enhanced version of diffusion model, termed as InstaGen, can serve as a data synthesizer, to enhance object detectors by training on its generated samples, demonstrating superior performance over existing state-of-the-art methods in open-vocabulary ($+4.5$ AP) and data-sparse ($+1.2\sim 5.2$ AP) scenarios.
@@ 3.3.2 Instance Grounding on Novel Categories
Training triplets of novel categories. We design the text prompts of novel categories, e.g., ‘a photograph of [novel category1] and [novel category2]’, and pass them through our proposed image synthesizer, to generate the visual features. To acquire the corresponding bounding-boxes for novel categories, we propose a self-training scheme that takes the above grounding head as the student, and apply a mean teacher (an exponential moving average (EMA) of the student model) to create pseudo labels for update. In contrast to the widely adopted self-training scheme that takes the image as input, the student and teacher in our case only take the visual features as input, thus cannot apply data augmentation as for images. Instead, we insert dropout module after each feature enhancer layer and decoder layer in the student. During training, we run inference (without dropout module) with teacher model on the visual features to produce bounding boxes, and then use a score threshold $\beta$ to filter out those with low confidence, and use the remaining training triplets $(\mathcal{F}_{i},\hat{b}_{i},y_{i}^{\text{novel}})$ to train the student, i.e., grounding head.
@@ 4.1 Implementation details
Subsequently, we train the instance grounding module with these training triplets for 6 epochs, with a batch size of 64. In the 6th epoch, we transfer the weights from the student model to the teacher model, and proceed to train the student for an additional 6 epochs. During this training, the student receives supervised training on the base categories and engages in self-training on novel categories, and the teacher model is updated using exponential moving average (EMA) with a momentum of 0.999. The initial learning rate is set to 1e-4 and is subsequently reduced by a factor of 10 at the 11-th epoch, and the score thresholds $\alpha$ and $\beta$ are set to 0.8 and 0.4, respectively.
@@ 4.2 Open-vocabulary object detection
Experimental setup. Following the previous works [5, 39], we conduct experiments on the open-vocabulary COCO benchmark, where 48 classes are treated as base categories, and 17 classes as the novel categories. To train the grounding head, we employ 1000 synthetic images per category per training epoch. While for training the object detector, we use 3000 synthetic images per category, along with the original real dataset for base categories. The object detector is trained with input size of $800\times 800$ and scale jitter. The performance is measured by COCO Average Precision at an Intersection over Union of 0.5 (AP50).
@@ 4.3 Data-sparse object detection
Experimental setup. Here, we evaluate the effectiveness of synthetic dataset in data-spare scenario, by varying the amount of real data. We randomly select subsets comprising 10%, 25%, 50%, 75% and 100% of the COCO training set, this covers all COCO categories. These subsets are used to fine-tune stable diffusion model for constructing image synthesizer, and train a Mask R-CNN for generating oracle ground-truth bounding boxes in synthetic images. We employ 1000 synthetic images per category to train a Faster R-CNN in conjunction with the corresponding COCO subset. The performance is measured by Average Precision [20].
@@ 4.4 Cross-dataset object detection
Experimental setup. In this section, we assess the effectiveness of synthetic data on a more challenging task, namely cross-dataset object detection. Following [39], we evaluate the COCO-trained model on two unseen datasets: Object365 [30] and LVIS [10]. Specifically, we consider the 48 classes in the open-vocabulary COCO benchmark as the source dataset, while Object365 (with 365 classes) and LVIS (with 1203 classes) serve as the target dataset. When training the instance grounding module, we acquire 1000 synthetic images for base categories from the source dataset, and 100 synthetic images for the category from the target dataset at each training iteration. In the case of training the object detector, we employ 500 synthetic images per category from the target dataset for each training iteration. The detector is trained with input size of $1024\times 1024$ and scale jitter [39].
@@ 6 Conclusion
This paper proposes a dataset synthesis pipeline, termed as InstaGen, that enables to generate images with object bounding boxes for arbitrary categories, acting as a free source for constructing large-scale synthetic dataset to train object detector. We have conducted thorough experiments to show the effectiveness of training on synthetic data, on improving detection performance, or expanding the number of detection categories. Significant improvements have been shown in various detection scenarios, including open-vocabulary ($+4.5$ AP) and data-sparse ($+1.2\sim 5.2$ AP) detection.
@@
7 Qualitative Results
We show more qualitative results generated by our InstaGen in Figure 5. Without any manual annotations, InstaGen can generate  high-quality  images  with  object  bounding-boxes
Figure 5: Qualitative results generated by our InstaGen. The bounding-boxes with green denote the objects from base categories, while the ones with red denote the objects from novel categories.
of novel categories. In Figure 6, we further show the qualitative results predicted by the Faster R-CNN trained with the synthetic images form InstaGen on COCO validation set. The detector can now accurately localize and recognize the objects from novel categories.
Figure 6: Qualitative results from our Faster R-CNN trained with the synthetic images from InstaGen on COCO validation set. The bounding-boxes with green denote the objects from base categories, while the ones with red denote the objects from novel categories.

### 08 Apr 2024 v3

@@ Abstract
In this paper, we present a novel paradigm to enhance the ability of object detector, e.g., expanding categories or improving detection performance, by training on synthetic dataset generated from diffusion models. Specifically, we integrate an instance-level grounding head into a pre-trained, generative diffusion model, to augment it with the ability of localising instances in the generated images. The grounding head is trained to align the text embedding of category names with the regional visual feature of the diffusion model, using supervision from an off-the-shelf object detector, and a novel self-training scheme on (novel) categories not covered by the detector. We conduct thorough experiments to show that, this enhanced version of diffusion model, termed as InstaGen, can serve as a data synthesizer, to enhance object detectors by training on its generated samples, demonstrating superior performance over existing state-of-the-art methods in open-vocabulary ($+4.5$ AP) and data-sparse ($+1.2\sim 5.2$ AP) scenarios.
@@ 3.3.2 Instance Grounding on Novel Categories
Training triplets of novel categories. We design the text prompts of novel categories, e.g., ‘a photograph of [novel category1] and [novel category2]’, and pass them through our proposed image synthesizer, to generate the visual features. To acquire the corresponding bounding-boxes for novel categories, we propose a self-training scheme that takes the above grounding head as the student, and apply a mean teacher (an exponential moving average (EMA) of the student model) to create pseudo labels for update. In contrast to the widely adopted self-training scheme that takes the image as input, the student and teacher in our case only take the visual features as input, thus cannot apply data augmentation as for images. Instead, we insert dropout module within each feature enhancer layer and decoder layer in the student. During training, we run inference (without dropout module) with teacher model on the visual features to produce bounding boxes, and then use a score threshold $\beta$ to filter out those with low confidence, and use the remaining training triplets $(\mathcal{F}_{i},\hat{b}_{i},y_{i}^{\text{novel}})$ to train the student, i.e., grounding head.
@@ 4.1 Implementation details
Subsequently, we train the instance grounding module with these training triplets for 6 epochs, with a batch size of 32. In the 6th epoch, we transfer the weights from the student model to the teacher model, and proceed to train the student for an additional 6 epochs. During this training, the student receives supervised training on the base categories and engages in self-training on novel categories, and the teacher model is updated using exponential moving average (EMA) with a momentum of 0.999. The initial learning rate is set to 1e-4 and is subsequently reduced by a factor of 10 at the 11th epoch, and the score thresholds $\alpha$ and $\beta$ are set to 0.8 and 0.4, respectively.
@@ 4.2 Open-vocabulary object detection
Experimental setup. Following the previous works [5, 39], we conduct experiments on the open-vocabulary COCO benchmark, where 48 classes are treated as base categories, and 17 classes as the novel categories. More results for LVIS can be found in the supplementary material. To train the grounding head, we employ 1250 synthetic images per category per training epoch. While for training the object detector, we use 3000 synthetic images per category, along with the original real dataset for base categories. The object detector is trained with input size of $800\times 800$ and scale jitter. The performance is measured by COCO Average Precision at an Intersection over Union of 0.5 (AP50).
@@ 4.3 Data-sparse object detection
Experimental setup. Here, we evaluate the effectiveness of synthetic dataset in data-spare scenario, by varying the amount of real data. We randomly select subsets comprising 10%, 25%, 50%, 75% and 100% of the COCO training set, this covers all COCO categories. These subsets are used to fine-tune stable diffusion model for constructing image synthesizer, and train a Mask R-CNN for generating oracle ground-truth bounding boxes in synthetic images. We employ 1250 synthetic images per category to train a Faster R-CNN in conjunction with the corresponding COCO subset. The performance is measured by Average Precision [20].
@@ 4.4 Cross-dataset object detection
Experimental setup. In this section, we assess the effectiveness of synthetic data on a more challenging task, namely cross-dataset object detection. Following [39], we evaluate the COCO-trained model on two unseen datasets: Object365 [30] and LVIS [10]. Specifically, we consider the 48 classes in the open-vocabulary COCO benchmark as the source dataset, while Object365 (with 365 classes) and LVIS (with 1203 classes) serve as the target dataset. When training the instance grounding module, we acquire 1250 synthetic images for base categories from the source dataset, and 100 synthetic images for the category from the target dataset at each training iteration. In the case of training the object detector, we employ 500 synthetic images per category from the target dataset for each training iteration. The detector is trained with input size of $1024\times 1024$ and scale jitter [39].
@@ 6 Conclusion
This paper proposes a dataset synthesis pipeline, termed as InstaGen, that enables to generate images with object bounding boxes for arbitrary categories, acting as a annotation-free approach for constructing large-scale synthetic dataset to train object detector. We have conducted thorough experiments to show the effectiveness of training on synthetic data, on improving detection performance, or expanding the number of detection categories. Significant improvements have been shown in various detection scenarios, including open-vocabulary ($+4.5$ AP) and data-sparse ($+1.2\sim 5.2$ AP) detection.
@@
In this supplementary document, we present the experimental results of the LVIS-OVD benchmark in Section S1. Additionally, we perform an ablation study to evaluate the coupling between the diffusion model and the grounding head in Section S2. Furthermore, we evaluate the quality of the pseudo-labels generated by the grounding head in Section S3. Lastly, we provide more qualitative results in Section S4.
S1 Open-vocabulary setting on LVIS
[…]
Similar to PromptDet [5], we enhance the prompt template by incorporating a more detailed description to mitigate lexical ambiguity, particularly for the rare classes in LVIS. It should be noted that the description can be easily extracted from the metadata of the dataset. Consequently, the text prompt for the selected categories is generated as follows: ‘a photograph of [category1 name] ([category1 description]) and [category2 name] ([category2 description])’. During the training of the grounding head, we utilize 500 synthetic images per category per training epoch. In addition, for the training of the object detector, we employ 250 synthetic images per category per training epoch and conduct 24 epochs of training.
Comparison to SOTA. We conduct a comparison with the existing CLIP-based open-vocabulary object detectors using the Mask-RCNN model with ResNet-50, as shown in Table S2. The results indicate that our detector, trained on synthetic dataset from InstaGen, achieves comparable or improved performance over existing CLIP-based methods.
S2 Tight coupling vs. Loose coupling
To generate high-quality bounding-boxes for the synthetic images, we have designed a tight coupling between the diffusion model and the instance-level grounding head, namely, the grounding head predicts the bounding-boxes based on the SDM’s internal representation. To demonstrate the effectiveness of the tight coupling design, we compare it with a loose coupling design. For the latter, we train an open-vocabulary detector (i.e. ResNet-101 + instance level grounding head) on the synthetic images with base categories, and generate pseudo-labels for novel categories. When training detectors on such synthetic dataset, it gives 31.9 AP on novel categories on the COCO-OVD benchmark, 10.4 AP lower than tight coupling, showing the benefits of rich semantic and positional information encoded in SDM’s visual features.
S3 Quality of Pseudo-labels
Here we evaluate the quality of the pseudo-labels generated by the proposed grounding head. We adopt two metrics to assess their quality: (i) Detector AP and (ii) Precision and Recall. For Detector AP, we leverage the pre-trained Mask-RCNN model on the COCO dataset to generate ground truths (GTs) for the synthetic images, and then compute the AP of the pseudo labels derived from the teacher model. In the case of Precision and Recall, we randomly select and annotate 200 synthetic images, then calculate the precision and recall of their pseudo-labels. As shown in Table S1, after self-training on novel categories, the quality of the pseudo-labels can be significantly improved in terms of Detector AP (70.2$\rightarrow$79.7%), Precision (87.9$\rightarrow$89.1%) and Recall (68.3$\rightarrow$90.0%).
$\mathcal{L}_{\text{base}}$ $\mathcal{L}_{\text{novel}}$ Detector AP Precision Recall
70.2 87.9 68.3
79.7 89.1 90.0
Table S1: The quality of the pseudo-labels.
S4 Qualitative Results
We show more qualitative results generated by our InstaGen in Figure S1. Without any manual annotations, InstaGen can generate high-quality images with object bounding-boxes of novel categories. In Figure S2, we further show the qualitative results predicted by the Faster R-CNN trained with the synthetic images form InstaGen on COCO validation set. The detector can now accurately localize and recognize the objects from novel categories.
[…]
ViLD-ens. [9] CLIP Mask R-CNN R50 1024$\times$1024 25.5 24.6 30.3 16.6
Detic [41] CLIP Mask R-CNN R50 1024$\times$1024 26.8 26.3 31.6 17.8
F-VLM [16] CLIP Mask R-CNN R50 1024$\times$1024 24.2 - - 18.6
PromptDet [5] CLIP Mask R-CNN R50 800$\times$800 21.4 18.5 25.8 19.0
DetPro [2] CLIP Mask R-CNN R50 800$\times$800 25.9 25.6 28.9 19.8
BARON [34] CLIP Mask R-CNN R50 800$\times$800 25.1 24.4 28.9 18.0
BARON [${}^{{\dagger}}$ CLIP Mask R-CNN R50 800$\times$800 27.6 27.6 29.8 22.6
InstaGen Stable Diffusion Mask R-CNN R50 800$\times$800 23.0 20.6 27.1 20.3
Table S2: Results on open-vocabulary LVIS benchmark. ${}^{{\dagger}}$ indicates using ensembling strategy for classification scores and learned prompts for the category’s names.
Figure S1: Qualitative results generated by our InstaGen. The bounding-boxes with green denote the objects from base categories, while the ones with red denote the objects from novel categories.
Figure S2: Qualitative results from our Faster R-CNN trained with the synthetic images from InstaGen on COCO validation set. The bounding-boxes with green denote the objects from base categories, while the ones with red denote the objects from novel categories.